OpenStudy (anonymous):

x squared + 18x -65=0 solve by factoring

OpenStudy (anonymous):

This polynomial can't be factored.

OpenStudy (anonymous):

how can't it be factored?

OpenStudy (accessdenied):

Are you sure it is not: x squared + 18x (+) 65 = 0 ? The plus sign would make it able to be factored.

OpenStudy (anonymous):

no the problem is x squared + 18x -65= 0

OpenStudy (accessdenied):

Strange, because then it is not factorable. -65 = -1 * 13*5 Your combinations are: 1, -65 -1, 65 -13, 5 13, -5. Note that one HAS to be negative because the end result is multiplying to obtain a negative: -65. None of these will add up to obtain +18 as an answer.

OpenStudy (accessdenied):

But if the problem were: x squared + 18x + 65 <-- with a plus sign there, 65 = 13 * 5 13, 5 do add up to 18. \( x^2 + 18x \color{green}+ 65 \color{green}= (x + 13)(x + 5) \) That wasn't the original problem, though. \( x^2 + 18x \color{red}- 65 \color{red}\ne (x + 13)(x + 5) \)

OpenStudy (anonymous):

maybe a typo

OpenStudy (accessdenied):

Ya. If this is a written assignment, I'd probably write "does not factor" and perhaps addend that it can be made factorable by that change. Not sure if in the context, there are problems that are supposed to be trick questions though. :)